OpenStudy (anonymous):

What is the percent of increase from 50 to 90? A. 40% B. 44% C. 56% D. 80%

OpenStudy (butterflydreamer):

find out what's the increase from 50 to 90. ( 90 - 50) = ?

OpenStudy (anonymous):

hmmmmm IM inclined to give a diff answer

OpenStudy (anonymous):

40

OpenStudy (vianne):

I'm pretty sure it's 80%..

OpenStudy (butterflydreamer):

Yep so now we need to find out what 40/50 = ?

OpenStudy (vianne):

50 to 90 90 - 50 = 40 40/50 = 80% increase

OpenStudy (butterflydreamer):

then multiply by 100 to get it as a percentage :)

OpenStudy (anonymous):

Yesssss

OpenStudy (anonymous):

80

OpenStudy (butterflydreamer):

that's correct ^_^

OpenStudy (anonymous):

Who do i medal?

OpenStudy (vianne):

Whoever you want. :) I think Butterfly helped the most than me. She guided you! ^-^

OpenStudy (anonymous):

Agreed ^

OpenStudy (anonymous):

OKay :)

OpenStudy (butterflydreamer):

:O you guys just increased my problem solving level :") <3 ty

OpenStudy (vianne):

Lol! Good Job.

OpenStudy (anonymous):

Np lol